### v2

@@ Appendix D Single-Action Auctions (Section 5) Appendix
The following definition for an extensive form auction is taken [^37] We modify the definition to remove notation we do not use and to make it specific to auctions. from Li (2017):
Definition O.1.
@@
      There exists an initial history $h_{\emptyset}=h$ such that there does not exists $h^{\prime}$ where $h^{\prime}\prec h$.
@@
    $\mathcal{A}:H\backslash h_{\emptyset}\to A$ maps histories to the most recent action taken to reach it. In addition:
@@
Example O.2.
Consider the sealed-bid, first-price (pay-as-bid) optimal auction with equilibrium bids $b^{1}$. The naïve implementation of allocation and transfer rule $(\tilde{x}^{*},\tilde{t}^{1})$ in a public auction would be to query each bidder sequentially on what her value is and then have the payment rule be $\tilde{t}^{1}$, but $\tilde{t}^{1}$ is not the direct transfer rule of any equilibrium of this game. Indeed, consider the last bidder who takes a move, and label that bidder $N$. If $v_{N}>\max_{i<N}\left\{b_{i}\right\}$, then the only possible equilibrium bid—and therefore the transfer function—is $\max_{i<N}\left\{b_{i}\right\}$.
@@
We can now write bidder $i$’s observations in the game as $((m_{i}^{k},r_{i}^{k})_{k=1}^{\tau_{i}},\omega_{i})$ where $\tau_{i}$ is the number of observations $i$ has and $\omega_{i}$ is the information partition over outcomes that $i$ observes. Let $o_{i}(\sigma_{0},\sigma,v)$ be $i$’s observation when the auctioneer plays $\sigma_{0}$, the bidders play $\sigma$, and the type profile is $v$.
Definition O.4 (Akbarpour and Li (2020)).
Let $\sigma_{0}^{G}$ be the rule-following auctioneer strategy. Formally, $\sigma_{0}^{G}$ is defined by the following algorithm: Initialize $\hat{h}:=h_{\emptyset}$. At each step, if $\hat{h}\in z$, end the game and choose outcome $(x,t)(\hat{h})$. Else, contact agent $P(\hat{h})$ and send message $m=\mathcal{I}_{P(\hat{h})}$ such that $P(\hat{h})\in\mathcal{I}_{P(\hat{h})}$. Upon receiving reply $r$, update $\hat{h}$ to $h$ such that $h\in\operatorname{succ}(\hat{h})$ and $\mathcal{A}(h)=r$, then iterate.
Definition O.5 (Akbarpour and Li (2020), Definition 3).
Given some promised strategy profile $(\sigma_{0},\sigma)$, auctioneer strategy $\hat{\sigma}_{0}$ is safe if, for all agents $i\in B$ and all type profiles $v$, there exists $v^{\prime}_{-i}$ such that $o_{i}(\hat{\sigma}_{0},\sigma,v)=o_{i}(\sigma_{0},\sigma,(v_{i},v^{\prime}_{-i}))$. We denote by $\Sigma^{*}_{0}(\sigma_{0},\sigma)$ the set of safe strategies.
Definition O.6 (Akbarpour and Li (2020), Definition 4).
@@
$\sigma^{G}_{0}\in\operatorname*{argmax}_{\sigma_{0}\in\Sigma^{*}_{0}(\sigma_{0}^{G},\sigma)}\left\{\mathds{E}_{v}\left[\sum_{i\in B}t_{i}(\sigma_{0},\sigma,v)\right]\right\}$
@@
Strong Shill-Proofness $\to$ Credibility. We prove the contrapositive: Suppose $(G,\sigma)$ is not credible. Let $\hat{\sigma}_{0}\in\Sigma^{*}_{0}(\sigma_{0}^{G},\sigma)$ be a profitable and safe deviation by the auctioneer. By Definition O.5, there exists $v$ and $\left\{v^{\prime}_{-i}\right\}_{i}$ such that $o_{i}(\hat{\sigma}_{0},\sigma,v)=o_{i}(\sigma_{0},\sigma,(v_{i},v^{\prime}_{-i}))$ for all $i$. By Lemma A.2, only one bidder $i^{*}$ has $t_{i^{*}}(\sigma(v_{i^{*}},v^{\prime}_{-i^{*}}))\neq 0$ and so $t_{i^{*}}(\sigma(v_{i^{*}},v^{\prime}_{-i^{*}}))>t_{i^{*}}(\sigma(v))$ because $\hat{\sigma}_{0}$ is profitable. Then, let $R=\left\{i^{*}\right\}$, and by ex-post monotonicity,
$t_{i^{*}}(\sigma(v_{i^{*}},v^{\prime}_{-i^{*}}))>t_{i^{*}}(\sigma(v))\geq t_{i^{*}}(\sigma(v_{i^{*}},0)),$
@@
Credibility $\to$ Weak Shill-Proofness. We prove the contrapositive: Suppose $(G,\sigma)$ is not weakly shill-proof. Let $\hat{\sigma}\in\Sigma_{S}$ be a profitable shilling strategy. Then, by Definition A.1, for all $R$, there exists $v,v^{\prime}$ such that $\hat{\sigma}(v;R)=\sigma(v;B)$. Consider the following reporting strategy $\hat{\sigma}_{0}$: for all $i\in R$, report play as if $i\in R$ is following $\hat{\sigma}$; and for all $i\notin R$, report in the rule-following manner. This strategy is safe because $\hat{\sigma}\in\Sigma_{S}$. To see that it is profitable compared to $\sigma_{0}^{G}$, consider what happens when the winning bidder $i$ is or is not in $R$ [^38] Note that if no one has value above the optimal reserve, there will be no winner under any safe strategy, so let us only consider the case where the good is allocated. Conditional on $i\in R$ winning, $\hat{\sigma}_{0}$ increases expected revenue because $\hat{\sigma}$ is a profitable shill bidding strategy. Conditional on $i\notin R$ winning, shill bidding would have led 00 revenue for the seller and, by Lemma A.2, $\hat{\sigma}_{0}$ must have non-negative revenue. Thus, our described reporting strategy is a profitable, safe strategy and therefore the auction is not credible. ∎
O.2 Generalizing Credibility in the Single-Action Case
Definition O.7.
Fix a single-action, optimal auction with exogenous signal $\psi$ and a set of real bidders $R$. The set of safe deviations to report to $i\in B$ is
$\mathcal{A}^{\psi}_{i}$ $(v_{j\leq i})=$
$\left\{a:\exists\tilde{v}_{-i}\text{ such that }\left[j<i,v_{j}\neq 0\implies\tilde{v}_{j}=v_{j}\right]\text{ and }a=\left(\sigma_{i}(v_{i};\psi_{i}(\tilde{v}_{j<i})),\sigma_{-i}\left(v_{i},\tilde{v}_{-i};B\right)\right)\right\}.$
@@
$\mathcal{A}^{\psi}(v)=\left\{\left\{a^{\leadsto i}\right\}:a^{\leadsto i}\in\mathcal{A}^{\psi}_{i}(v_{j\leq i})\text{ and }\sum_{i=1}^{N}x_{i}\left(a^{\leadsto i}\right)\leq 1\right\}.$
Definition O.7 allows the auctioneer to report any value she chooses when a bidder’s declared valuation is 00. Note that if we take the canonical setting where there are no exogenous signals, the above assumption is without loss.
Definition O.8.
A single-action, optimal auction is $\psi$-credible if for all $v$ and $\left\{a^{\leadsto i}\right\}\in\mathcal{A}^{\psi}(v)$, we have
$\sum_{i}t_{i}(a^{\leadsto i})\leq\sum_{i}t_{i}(\sigma(v;B)).$
Lemma O.9.
For a single-action, optimal auction, define the augmented (direct) inverse $\widecheck{\psi}_{i}^{-1}$ as $\widecheck{\psi}_{i}^{-1}(v)=\left\{0\right\}\cup\psi_{i}^{-1}(v_{j<i})$. Then for
$\mathcal{V}^{\psi}(v)=\left\{\left\{v^{\leadsto i}\right\}:v^{\leadsto i}\in\widecheck{\psi}^{-1}_{i}(v),\sum_{i}\tilde{x}^{*}_{i}(v^{\leadsto i})\leq 1\right\},$
the auction is credible if and only if for all $v$, and $\left\{v^{\leadsto i}\right\}\in\mathcal{V}^{\psi}(v)$,
$\sum_{i}\tilde{t}_{i}(v^{\leadsto i})\leq\sum_{i}\tilde{t}_{i}(v).$
@@
Apply Lemma D.1, specifically the unique mapping between $(\tilde{x}^{*},\tilde{t})$ and $(x,t)$ to Definitions O.7 and O.6 to see that the lemma holds. ∎
Lemma O.10.
Suppose a single-action, optimal auction is weakly shill-proof, but not strongly shill-proof. Then, there exist $R$, $v_{R}$, and $v_{-R}$ such that
$\sum_{k\in R}\tilde{t}_{k}(v_{R},v_{-R})>\sum_{k\in R}\tilde{t}_{k}(v_{R},0).$ (15)
@@
Suppose that $(G,\sigma)$ is weakly shill-proof, but not strongly shill-proof. Because $(G,\sigma)$ is weakly shill-proof, for all $v$ and $R,R^{\prime}$, we can define $\hat{\sigma}(v)\equiv\sigma(v;R)=\sigma(v;R^{\prime})$. Since $(G,\sigma)$ is not strongly shill-proof, $\hat{\sigma}$ must not be an ex-post strategy for the shill bidders. So, for some realization of $R$ and $v_{R}$ there exists a profitable deviation for the shill bidders; examining the set of possible deviations $\Sigma_{S}$ in Definition A.1, we see that any profitable deviating actions induces a profitable misreport $v_{-R}$ in the direct mechanism for some $R$ and $v_{R}$; proving Equation 15 can be satisfied. ∎
@@
Weak Shill-Proofness $\to$ $(\psi=\mathrm{Id})$-Credibility. Suppose the auction is not $(\psi=\mathrm{Id})$-credible. Then, combining Lemma O.9 with the ex-post IR condition, there exist $v,\left\{v^{\leadsto i}\right\}\in\mathcal{V}^{\psi}(v)$ and $k^{*}$ such that $\tilde{t}_{k^{*}}(v^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(v)$. Applying the winner-paying property, it is the case that for all $j\neq k^{*}$, $\tilde{t}_{j}(v^{\leadsto k^{*}})=0$. Since $\psi=\mathrm{Id}$, for all $j\leq k$, it is the case that $v^{\leadsto k^{*}}_{j}=v_{j}$ or $v_{j}=0$. Let $R=\left\{1,\ldots,k^{*}\right\}$. Then,
$\sum_{i\in R}\tilde{t}_{i}(v^{\leadsto k^{*}})$ $=\tilde{t}_{k^{*}}(v^{\leadsto k^{*}})$
$\geq\tilde{t}_{k^{*}}\left(v_{1},\ldots,v_{k^{*}},v^{\leadsto k^{*}}_{k^{*}+1},\ldots,v^{\leadsto k^{*}}_{N}\right)$
$\geq\tilde{t}_{k^{*}}\left(v_{1},\ldots,v_{k^{*}},0,\ldots,0\right)$
$=\sum_{i\in R}\tilde{t}_{i}(v_{R},0).$
Thus, we can apply Lemma D.2 to conclude that the auction is not weakly shill-proof.
$\psi$-Credibility $\to$ Weak Shill-Proofness. Suppose the auction is $\psi$-credible. Towards contradiction, suppose the auction is not weakly shill-proof. So, there exists $R$ and $v$ such that $\sigma(v;R)\neq\sigma(v;B)$. In particular, this means that shill bidders have, in expectation, a profitable deviation relative to acting as real bidders with valuation 00. If this is true in expectation, there must then exist $v=(v_{R},0)$ and $\tilde{v}_{-R}$ such that
$\sum_{i\in R}\tilde{t}_{i}((v_{R},\tilde{v}_{-R}))>\sum_{i\in R}\tilde{t}_{i}((v_{R},0)).$

### v3

@@ Appendix E Supplemental Appendix
The following definition for an extensive-form auction is taken [^34] We modify the definition to remove notation we do not use and to make it specific to auctions. from Li (2017):
Definition E.1.
@@
      There exists an initial history $h_{\emptyset}$ such that there does not exist $h^{\prime}$ where $h^{\prime}\prec h_{\emptyset}$.
@@
[…]
@@
In order to define an information order, we will use the notation that when the current set of possible values is $\Theta$, a player $i$’s knowledge of what values are possible is $\mathcal{I}^{i}_{\Theta}$.
Definition E.2.
A menu rule $\mu^{\prime}$ is more informative than $\mu$, $\mu^{\prime}\succeq_{\mathrm{Info}}\mu$, if $\mu(\cdot,\cdot)=\mu^{\prime}(\cdot,\cdot)$ for all game states $\Theta$ and all possible information sets $\mathcal{I},\mathcal{I}^{\prime}$, when $\mathcal{I}\subseteq\mathcal{I}^{\prime}$, it is the case that
$\mathds{P}\left[\mathcal{I}_{\Theta}=\mathcal{I}\mid\mu^{\prime}\right]-\mathds{P}\left[\mathcal{I}_{\Theta}=\mathcal{I}^{\prime}\mid\mu^{\prime}\right]\geq\mathds{P}\left[\mathcal{I}_{\Theta}=\mathcal{I}\mid\mu\right]-\mathds{P}\left[\mathcal{I}_{\Theta}=\mathcal{I}^{\prime}\mid\mu\right].$
Proposition E.3.
Consider any affiliated type distribution $F$ and value function $v$ satisfying Assumption 5.1. Suppose $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu,\xi_{0},v,F),\mu,\xi_{0})$ is orderly and strongly shill-proof. Then, for any $\mu^{\prime}$ such that $\mu\succeq_{\mathrm{Info}}\mu^{\prime}$, it is the case that $(\tilde{x},\tilde{t}^{*}(\tilde{x},\mu^{\prime},\xi_{0},v,F),\mu^{\prime},\xi_{0})$ is strongly shill-proof.
Proof.
Towards contrapositive, suppose that the auction is not strongly shill-proof under $\mu$ and consider some $\mu^{\prime}\succeq_{\mathrm{Info}}\mu$. There exists some realization of $\theta$ under $F$ such that a shill bidder $i$ deviates at $\Theta$. Since $F$ is affiliated and $\tilde{x}$, $\xi_{0}$, and $\mu$ are all fixed, we can apply Lemmatta D.1, 5.5 and B.3 to conclude that the only way for expected revenue under strong shill-proofness to change is for ex-interim expected values for the winner $j\neq i$ to change. There can be no change in allocation since $j$ has the highest value among the real bidders and so since the allocation is orderly, any change in the allocation would result in a shill bidder winning the item. Recall that the winning bidder conditions her valuation on the realization that she wins, i.e., she conditions on the assumption that $(\theta_{j},j)\rhd(\theta_{k},k)$ for all $k\neq j$. Thus, it is without loss with regards to how the winning bidder $j$ will estimate her valuation to assume that $\mathcal{I}_{\Theta}$ is such that for all $y\in\mathcal{I}_{\Theta}$ and $k\neq j$, $(\theta_{j},j)\rhd(y_{k},k)$. There are now two cases to consider. In the first case, consider that under $\mu^{\prime}$, the information sets in the support are a super-set of those in the support under $\mu$. We then condition on the event of the same realization of the information set that leads to a profitable deviation to find that the auction is not strongly shill-proof under $\mu^{\prime}$. In the second case, we do not assume that the support under $\mu$ contains elements that $\mu^{\prime}$ does not. It may be that the realization of the information set for the winner’s last move is not in the support of $\mu$. However, because $\mu^{\prime}\succeq_{\mathrm{Info}}\mu$, we can know that there is some realization of the information structure such that 0 is not contained in it. This is because we know that there exists an information set that is strictly smaller than that realization and by Bayes plausibility, the complement must also be in it. Further, a shill bidding deviation means selecting a partition which does not include 0 and by Lemma A.4 Condition 2a, we know that said partition does not intersect the partition containing 0. Since the allocation is fixed, one of the partitions must increase transfers from the winner in order to have the average transfer under every realization of the information set average out to the transfer under the coarser information set. We then select the highest transfer among these realizations to find a profitable deviation and complete the proof. ∎
Example E.4.
Consider the sealed-bid, first-price (pay-as-bid) optimal auction with equilibrium bids $b^{1}$ in the regular, IPV environment. The naïve implementation of allocation and transfer rule $(\tilde{x}^{*},\tilde{t}^{1})$ in a public auction would be to query each bidder sequentially on what her value is and then have the payment rule be $\tilde{t}^{1}$, but $\tilde{t}^{1}$ is not the direct transfer rule of any equilibrium of this game. Indeed, consider the last bidder who takes a move, and label that bidder $N$. If $\theta_{N}>\max_{i<N}\left\{b_{i}\right\}$, then the only possible equilibrium bid—and therefore the transfer function—is $\max_{i<N}\left\{b_{i}\right\}$.
@@ E.1 Credibility
We can now write bidder $i$’s observations in the game as $((m_{i}^{k},r_{i}^{k})_{k=1}^{\tau_{i}},\omega_{i})$ where $\tau_{i}$ is the number of observations $i$ has and $\omega_{i}$ is the information partition over outcomes that $i$ observes. Let $o_{i}(\sigma_{0},\sigma,\theta)$ be $i$’s observation when the auctioneer plays $\sigma_{0}$, the bidders play $\sigma$, and the type profile is $\theta$.
Definition E.6 (Akbarpour and Li (2020)).
[…]
Definition E.7 (Akbarpour and Li (2020), Definition 3).
Given some promised strategy profile $(\sigma_{0},\sigma)$, auctioneer strategy $\hat{\sigma}_{0}$ is safe if, for all agents $i\in B$ and all type profiles $\theta$, there exists $\theta^{\prime}_{-i}$ such that $o_{i}(\hat{\sigma}_{0},\sigma,\theta)=o_{i}(\sigma_{0},\sigma,(\theta_{i},\theta^{\prime}_{-i}))$. We denote by $\Sigma^{*}_{0}(\sigma_{0},\sigma)$ the set of safe strategies.
Definition E.8 (Akbarpour and Li (2020), Definition 4).
@@
$\sigma^{G}_{0}\in\operatorname*{argmax}_{\sigma_{0}\in\Sigma^{*}_{0}(\sigma_{0}^{G},\sigma)}\left\{\mathds{E}_{\theta}\left[\sum_{i\in B}t_{i}(\sigma_{0},\sigma,\theta)\right]\right\}$
@@
Strong Shill-Proofness $\to$ Credibility. We prove the contrapositive: Suppose $(G,\sigma)$ is not credible. Let $\hat{\sigma}_{0}\in\Sigma^{*}_{0}(\sigma_{0}^{G},\sigma)$ be a profitable and safe deviation by the auctioneer. By Definition E.7, there exists $\theta$ and $\left\{\theta^{\prime}_{-i}\right\}_{i}$ such that $o_{i}(\hat{\sigma}_{0},\sigma,\theta)=o_{i}(\sigma_{0},\sigma,(\theta_{i},\theta^{\prime}_{-i}))$ for all $i$. By Lemma B.1, only one bidder $i^{*}$ has $t_{i^{*}}(\sigma(\theta_{i^{*}},\theta^{\prime}_{-i^{*}}))\neq 0$ and so $t_{i^{*}}(\sigma(\theta_{i^{*}},\theta^{\prime}_{-i^{*}}))>t_{i^{*}}(\sigma(\theta))$ because $\hat{\sigma}_{0}$ is profitable. Then, let $R=\left\{i^{*}\right\}$, and by ex-post monotonicity,
$t_{i^{*}}(\sigma(\theta_{i^{*}},\theta^{\prime}_{-i^{*}}))>t_{i^{*}}(\sigma(\theta))\geq t_{i^{*}}(\sigma(\theta_{i^{*}},0)),$
@@
Credibility $\to$ Weak Shill-Proofness. We prove the contrapositive: Suppose $(G,\sigma)$ is not weakly shill-proof. Let $\hat{\sigma}\in\Sigma_{S}$ be a profitable shilling strategy. Then, by Definition A.1, for all $R$, there exists $\theta,\theta^{\prime}$ such that $\hat{\sigma}(\theta;R)=\sigma(\theta;B)$. Consider the following reporting strategy $\hat{\sigma}_{0}$: for all $i\in R$, report play as if $i\in R$ is following $\hat{\sigma}$; and for all $i\notin R$, report in the rule-following manner. This strategy is safe because $\hat{\sigma}\in\Sigma_{S}$. To see that it is profitable compared to $\sigma_{0}^{G}$, consider what happens when the winning bidder $i$ is or is not in RR. [^35] Note that if no one has value above the optimal reserve, there will be no winner under any safe strategy, so let us only consider the case where the good is allocated. Conditional on $i\in R$ winning, $\hat{\sigma}_{0}$ increases expected revenue because $\hat{\sigma}$ is a profitable shill bidding strategy. Conditional on $i\notin R$ winning, shill bidding would have led 0 revenue for the seller and, by Lemma B.1, $\hat{\sigma}_{0}$ must have non-negative revenue. Thus, our described reporting strategy is a profitable, safe strategy and therefore the auction is not credible. ∎
E.2 Generalizing Credibility in the Single-Action Case
Definition E.9.
Fix a single-action auction with exogenous signal $\psi$ and a set of real bidders $R$. The set of safe deviations to report to $i\in B$ is
$\mathcal{A}^{\psi}_{i}$ $(\theta_{j\leq i})=$
$\left\{a:\exists\tilde{\theta}_{-i}\text{ such that }\left[j<i,\theta_{j}\neq 0\implies\tilde{\theta}_{j}=\theta_{j}\right]\text{ and }a=\left(\sigma_{i}(\theta_{i};\psi_{i}(\tilde{\theta}_{j<i})),\sigma_{-i}\left(\theta_{i},\tilde{\theta}_{-i};B\right)\right)\right\}.$
@@ E.2 Generalizing Credibility in the Single-Action Case
$\mathcal{A}^{\psi}(\theta)=\left\{\left\{a^{\leadsto i}\right\}:a^{\leadsto i}\in\mathcal{A}^{\psi}_{i}(\theta_{j\leq i})\text{ and }\sum_{i=1}^{N}x_{i}\left(a^{\leadsto i}\right)\leq 1\right\}.$
Definition E.9 allows the auctioneer to report any value she chooses when a bidder’s declared valuation is 0. Note that if we take the canonical setting where there are no exogenous signals, the above assumption is without loss.
Definition E.10.
A single-action auction is $\psi$-credible if for all $\theta$ and $\left\{a^{\leadsto i}\right\}\in\mathcal{A}^{\psi}(\theta)$, we have
$\sum_{i}t_{i}(a^{\leadsto i})\leq\sum_{i}t_{i}(\sigma(\theta;B)).$
Lemma E.11.
For a single-action auction, define the augmented (direct) inverse $\widecheck{\psi}_{i}^{-1}$ as $\widecheck{\psi}_{i}^{-1}(\theta)=\left\{0\right\}\cup\psi_{i}^{-1}(\theta_{j<i})$. Then for
$\vartheta^{\psi}(\theta)=\left\{\left\{\theta^{\leadsto i}\right\}:\theta^{\leadsto i}\in\widecheck{\psi}^{-1}_{i}(\theta),\sum_{i}\tilde{x}_{i}(\theta^{\leadsto i})\leq 1\right\},$
the auction is credible if and only if for all $\theta$, and $\left\{\theta^{\leadsto i}\right\}\in\vartheta^{\psi}(\theta)$,
$\sum_{i}\tilde{t}_{i}(\theta^{\leadsto i})\leq\sum_{i}\tilde{t}_{i}(\theta).$
@@
[…]
Lemma E.12.
Suppose a single-action auction is weakly shill-proof, but not strongly shill-proof. Then, there exist $R$, $\theta_{R}$, and $\theta_{-R}$ such that
$\sum_{k\in R}\tilde{t}_{k}(\theta_{R},\theta_{-R})>\sum_{k\in R}\tilde{t}_{k}(\theta_{R},0).$ (17)
@@
Suppose that $(G,\sigma)$ is weakly shill-proof, but not strongly shill-proof. Because $(G,\sigma)$ is weakly shill-proof, for all $\theta$ and $R,R^{\prime}$, we can define $\hat{\sigma}(\theta)\equiv\sigma(\theta;R)=\sigma(\theta;R^{\prime})$. Since $(G,\sigma)$ is not strongly shill-proof, $\hat{\sigma}$ must not be an ex-post strategy for the shill bidders. So, for some realization of $R$ and $\theta_{R}$ there exists a profitable deviation for the shill bidders; examining the set of possible deviations $\Sigma_{S}$ in Definition A.1, we see that any profitable deviating actions induces a profitable misreport $\theta_{-R}$ in the direct mechanism for some $R$ and $\theta_{R}$; proving Equation 17 can be satisfied. ∎
Lemma E.13.
If a single-action auction is strongly shill-proof, then for all $R$, $i\notin R$, $\theta_{i}$, and $\theta_{-i}$, $\sum_{k\in R}\tilde{t}_{k}(\theta_{i},\theta_{-i})\leq\sum_{k\in R}\tilde{t}_{k}(0,\theta_{-i}).$
Proof.
Towards contradiction, suppose that there exists $R$, $i\notin R$, $\theta_{i}$, and $\theta_{-i}$ such that $\sum_{k\in R}\tilde{t}_{k}(\theta_{i},\theta_{-i})>\sum_{k\in R}\tilde{t}_{k}(0,\theta_{-i}).$ So in the direct game reporting 0 is not a dominant strategy for shill bidders. This implies, from Condition i of Lemma C.1, that there exists a deviation in the auction such that for some value vectors, the seller raises more revenue. Therefore, the auction is not strongly shill-proof. ∎
@@
Weak Shill-Proofness $\to$ $(\psi=\mathrm{Id})$-Credibility. Suppose the auction is not $(\psi=\mathrm{Id})$-credible. Then, combining Lemma E.11 with the ex-post IR condition, there exist $\theta,\left\{\theta^{\leadsto i}\right\}\in\vartheta^{\psi}(\theta)$ and $k^{*}$ such that $\tilde{t}_{k^{*}}(\theta^{\leadsto k^{*}})>\tilde{t}_{k^{*}}(\theta)$. Applying the definition of orderly and the winner-paying property, for all $j\neq k^{*}$, $\tilde{t}_{j}(\theta^{\leadsto k^{*}})=0$. Since $\psi=\mathrm{Id}$, for all $j\leq k$, it is the case that $\theta^{\leadsto k^{*}}_{j}=\theta_{j}$ or $\theta_{j}=0$. Let $R=\left\{1,\ldots,k^{*}\right\}$. Then,
$\sum_{i\in R}\tilde{t}_{i}(\theta^{\leadsto k^{*}})$ $=\tilde{t}_{k^{*}}(\theta^{\leadsto k^{*}})$
$\geq\tilde{t}_{k^{*}}\left(\theta_{1},\ldots,\theta_{k^{*}},\theta^{\leadsto k^{*}}_{k^{*}+1},\ldots,\theta^{\leadsto k^{*}}_{N}\right)$
$\geq\tilde{t}_{k^{*}}\left(\theta_{1},\ldots,\theta_{k^{*}},0,\ldots,0\right)$
$=\sum_{i\in R}\tilde{t}_{i}(\theta_{R},0).$
Thus, we can apply Lemma C.2 to conclude that the auction is not weakly shill-proof.
$\psi$-Credibility $\to$ Weak Shill-Proofness. Suppose the auction is $\psi$-credible. Towards contradiction, suppose the auction is not weakly shill-proof. So, there exists $R$ and $\theta$ such that $\sigma(\theta;R)\neq\sigma(\theta;B)$. In particular, this means that shill bidders have, in expectation, a profitable deviation relative to acting as real bidders with valuation 0. If this is true in expectation, there must then exist $\theta=(\theta_{R},0)$ and $\tilde{\theta}_{-R}$ such that
$\sum_{i\in R}\tilde{t}_{i}((\theta_{R},\tilde{\theta}_{-R}))>\sum_{i\in R}\tilde{t}_{i}((\theta_{R},0)).$